# Remote Auditing: Design-Based Tests of Randomization, Selection, and Missingness with Broadly Accessible Satellite Imagery

### v1

@@ 5 What Else Should Analysts Check in Remote Audits?
Structure-destroying placebos—such as random rotations of patches or within-region permutations of embeddings—help verify that the pipeline does not manufacture predictability. Complementary stress tests can inject stylized deviations that mimic realistic implementation failures (e.g., favoritism toward road-adjacent or administrative-boundary units) to assess whether the audit would detect such problems at application-relevant sample sizes. A simple simulation calibrated to the study provides a practical power check.
When the audit suggests potential issues, investigators typically consider complementary design or analysis steps:
  • Randomization flagged. Teams often tighten operational constraints or add stratification in a re-randomization, strengthen monitoring of assignment processes (seeds, logs, audit trails), and document the deviation together with pre-specified adaptations for inference (e.g., design-based variance, randomization inference, or sensitivity analysis). If re-randomization occurs, the audit is usually re-run and archived alongside the final assignment.
@@
  • Selection flagged. It can be useful to revisit the sampling frame and clarify the target-population estimand, consider reweighting or targeted enrollment to improve coverage of under-represented areas, and state limits to external validity transparently; transported or reweighted estimates may be offered as secondary analyses where appropriate.
@@
  • Missingness flagged. Common responses include additional follow-up in hard-to-reach locations, small adaptations to instruments or field protocols, and pre-specification of imputation or inverse-probability weighting (avoiding complete-case analyses when missingness is predictably related to $\boldsymbol{\phi}$); attrition bounds provide an additional robustness assessment.
Most of these adjustments are far cheaper and cleaner before enumeration begins. That is the practical advantage of an imagery-only audit: it surfaces potential problems early enough that designs can still be improved rather than merely documented.
6 Scope, Risks, and Implementation Guidance
Remote Audit in Context.
Researchers already exploit satellite imagery to measure outcomes, build covariates, and study heterogeneity (Jean et al., 2016; Yeh et al., 2020; Jerzak, Johansson and Daoud, 2023; Torres and Pugh, 2022). Imagery has also been used model-based to mitigate confounding when pre-treatment signals proxy latent conditions (Sanford, 2021; Burke et al., 2021). Our contribution is orthogonal to these uses. We deploy imagery in a design-based capacity: a conditional randomization test that asks whether realized assignment ($A_{i}$) is independent of pre-treatment embeddings ($\boldsymbol{\phi}_{i}$) under the registered mechanism ($\Omega$). Because the audit never adjusts outcomes, it avoids “bad control” pitfalls and mediator/collider concerns that attend post-treatment measurement in model-based pipelines (Angrist and Pischke, 2009; Pearl, 2009; Cinelli and Hazlett, 2020). Where model-based approaches typically require identification assumptions and sensitivity analysis (Rosenbaum and Rubin, 1983; Oster, 2019), our audit derives validity from the design itself: extremeness is judged against the experiment’s finite-sample reference generated by draws from ($\Omega$). In this sense, the paper reframes what imagery is for—not as a control set inside an outcome model, but as a ubiquitous, pre-existing source of design-stage information capable of certifying whether assignment behaves as if random.
What satellites can test—and what they cannot.
@@ 6 Scope, Risks, and Implementation Guidance
Visual signals reliably register elements of the built and natural environment—settlement structure, roads, roof materials, vegetation, hydrology, and some economic activity such as night lights (Henderson, Storeygard and Weil, 2012; Jean et al., 2016; Watmough et al., 2019). These are precisely the kinds of context that often co-move with accessibility, wealth, and administrative capacity. A rejection by our audit therefore indicates an image-aligned deviation from ($\Omega$); it does not identify mechanisms or actors, which require field investigation. Conversely, a non-rejection is not a certificate of perfection: many forces relevant to assignment (patronage networks, norms, internal procedures) leave weak or no satellite trace at available resolutions. The paper’s insistence on strictly pre-treatment imagery is central to this point. Using images captured after mobilization risks encoding mediators (construction, publicity) or conditioning on colliders (selection into measurement), with familiar causal consequences (Angrist and Pischke, 2009; Pearl, 2009). Our guidance is to commit ex ante to a pre-treatment window, to document sensor choice, compositing, and quality control, and to quantify geocoding uncertainty with patch sizes (s) that are robust to plausible error.
Principles for valid and reproducible audits.
The audit’s guarantees hinge on separating representation from inference and evaluating strictly out of sample. We therefore precompute embeddings—interpretable indices and/or off‑the‑shelf encoders—fixing checkpoints, layers, and normalizations ex ante; we share hashes and manifests so that ($\boldsymbol{\phi}$) can be reproduced byte-for-byte. We fit simple learners first, deploying more complicated models only as needed, and use repeated (K)-fold cross-fitting with seeds recorded. Clustered designs should leave whole clusters out of folds to mitigate spatial leakage. The statistic is defined as out‑of‑sample log‑likelihood improvement relative to the assignment probabilities implied by ($\Omega$), so that “no signal” maps transparently to zero. Because our test is randomization‑based, it naturally honors blocks and clusters during resampling, and its pp‑values are finite‑sample valid without parametric assumptions. These elements make the audit naturally preregistrable: the image window, representation, folds, statistic, and resampling rule can all be locked in advance, limiting the researcher’s degrees of freedom (Bouguen et al., 2019).
Multiplicity, computation, and decision rules.
Exploration across embeddings, resolutions, and patch sizes is often scientifically reasonable but inflates Type I error if not considered in analysis. We therefore recommend Westfall–Young max‑$T$ adjustment as a powerful, design‑compatible correction; simple conservative alternatives (Bonferroni, BH‑style control) are available when appropriate (Westfall and Young, 1993; Thissen, Steinberg and Kuang, 2002). Preregistered decision rules should specify thresholds (e.g., $\alpha$ = 0.05) and remediation steps—re‑randomization, tightening operational constraints, or proceeding with transparent documentation when re‑randomization is infeasible—treating the audit as a design‑stage gate akin to power analysis.
Interpretability, communication, and ethics.
Although the audit’s validity does not depend on interpretability, policy audiences benefit from understanding “what the model saw.” Post‑hoc summaries such as permutation importance or representative patch visualizations can aid communication, provided we separate explanation from inference. Finally, because satellite data may be granular, ethical use requires attention to geospatial privacy and local norms. Sharing processed features and manifests rather than raw tiles, when feasible, reduces re‑identification risk without compromising the design-based logic of the audit.
@@
In sum, the remote audit’s distinctive contribution is to repurpose broadly accessible pre‑treatment imagery for a genuinely design‑based question—did realized assignment conform to ($\Omega$)?—and to do so with a preregistrable, finite‑sample‑valid procedure that scales across stratified and clustered experiments. This positions remote audits as a low‑cost complement to conventional diagnostics: powerful when deviations align with visible context, transparent about their limits when they do not, and deployable early enough to improve designs rather than merely document them.
@@ 7 Conclusion
We develop and demonstrate a remote audit of randomization integrity that leverages only pre-treatment satellite imagery and a conditional randomization test. The audit is valid in finite samples, easily preregistered, and compatible with stratified and clustered designs. In Uganda’s Youth Opportunities Program (Blattman, Fiala and Martinez, 2014), it would have corroborated the reported lottery while flagging selection and missing-data risks that matter for interpretation and design; in a field experiment in Bangladesh, randomization integrity is itself questioned, in line with concerns raised by an independent team of investigators.
Remote audits are not a panacea. The remote audit detects only those design-context correlations that are captured by pre-treatment imagery. Some factors that influence assignment may be invisible from space or imperfectly proxied at available resolutions. Geocoding errors and cloud cover can reduce power. A non-rejection is not a proof of integrity, and a rejection does not uniquely identify the source of deviation. For selection and missingness extensions, the tests are descriptive rather than design-based. Despite these limits, the audit adds a transparent, low-cost, preregistrable safeguard where ground measurement is difficult.
As global image archives deepen and off-the-shelf vision models improve (Li et al., 2020; Dosovitskiy et al., 2020; Liu et al., 2021), the cost of running such audits is becoming vanishingly small. For experiments in challenging environments—where baseline enumeration is costly, schedules tight, and logistics uncertain—remote audits offer a practical early-warning system. We recommend incorporating them into preregistration and routine experimental quality assurance alongside conventional balance checks and process documentation (Bouguen et al., 2019). $\square$
@@
Proof. Condition on $(\boldsymbol{\phi},H)$. Under the null, $A\sim\Omega$ and $A^{(1)},\ldots,A^{(B)}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\Omega$ are exchangeable. Applying the fixed, measurable map $g(\cdot)$ to each assignment yields exchangeable statistics $\big(T,T^{(1)},\ldots,T^{(B)}\big)$. Therefore the rank of $T$ among these B+1B{+}1 values is uniformly distributed on $\{1,\ldots,B{+}1\}$ (with ties handled by the $\geq$ rule), which implies that $p$ is super-uniform and $\Pr(p\leq\alpha\mid\boldsymbol{\phi},H)\leq\alpha$. When $\Omega$ imposes block or cluster totals, exchangeability holds conditional on those totals, so the same argument applies. $\square$
Remark. Refitting the predictive model within each resample is what ensures that $g(\boldsymbol{\phi},\cdot,S)$ treats every draw from $\Omega$ symmetrically; reusing a learner trained only on the realized AA can break exchangeability. See Candes et al. (2018) and Hennessy et al. (2016) for more information.

### v2

@@ 6 Discussion
Scale and sensing choices also matter. Vary image patch sizes to cover plausible geocoding error and local spillovers, and, where multiple pre-treatment sensors exist, consider parallel analyses. Placebo tests that destroy structure, such as remote audits of synthetically allocated treatments, help verify that the pipeline does not manufacture predictability. Complementary stress tests can inject stylized deviations that mimic realistic implementation failures—favoritism toward road-adjacent or administrative-boundary units—to gauge whether the audit would detect such problems at application-relevant sample sizes; a simple, study-calibrated simulation can provide a practical power check.
When the audit suggests potential issues, responses are straightforward. If randomization is flagged, teams might tighten operational constraints or add stratification and re-randomize. They might also strengthen monitoring (seeds, implementation logs) and document deviations.
@@
If selection is flagged, revisit the sampling frame, clarify the target-population estimand, consider reweighting or targeted enrollment to improve coverage of under-represented areas, and state external validity limits (Mullinix et al., 2015); transported or reweighted estimates may be offered as secondary analyses where appropriate.
@@
If missingness is flagged, plan additional follow-up in hard-to-reach locations, make small adaptations to instruments or field protocols, and pre-specify imputation or inverse-probability weighting (avoiding complete-case analyses when missingness is predictably related to imagery); attrition bounds provide an additional robustness assessment. Most adjustments are cheaper and cleaner before enumeration begins—the practical advantage of an imagery-only audit is precisely that it surfaces potential problems early enough to improve designs rather than merely document them.
@@
What satellites can test—and what they cannot—is an open and important question. Visual signals reliably register elements of the built and natural environment—settlement structure, roads, roof materials, vegetation, hydrology, and some economic activity such as night lights (Henderson, Storeygard and Weil, 2012; Jean et al., 2016; Watmough et al., 2019). These features often co-move with accessibility, wealth, and administrative capacity. A rejection, therefore, indicates an image-aligned deviation from $(\Omega)$; it does not identify mechanisms or actors, which require field investigation. Conversely, a non-rejection is not a certificate of perfection: many forces relevant to assignment (patronage networks, norms, internal procedures) leave weak or no satellite trace at available resolutions. The insistence on strictly pre-treatment imagery is central. Using images captured after mobilization risks encoding mediators (construction, publicity) or conditioning on colliders (selection into measurement), with familiar causal consequences.
Although validity does not depend on interpretability, policy audiences benefit from understanding “what the model saw.” Here, comparing interpretable indices—vegetation, built-up, texture—with learned neural network encoders helps facilitate interpretation of how treatment and control differ. Also, post-hoc summaries such as feature importance or representative patch visualizations can aid communication, provided explanation is kept separate from inference.
@@
In sum, remote audits repurpose broadly accessible pre-treatment imagery to answer a genuinely design-based question—did realized assignment conform to the design? or there residual imbalance, even if the design was faithfully followed?—with a preregistrable, finite-sample-valid procedure that scales across stratified and clustered experiments. They offer a low-cost and fast complement to conventional diagnostics: powerful when deviations align with visible context and deployable early enough to improve designs rather than merely document them.
@@ 7 Conclusion
We develop and demonstrate a remote audit of randomization integrity that leverages only pre-treatment satellite imagery and a conditional randomization test. The audit is valid in finite samples, easily preregistered, and compatible with stratified and clustered designs. In Uganda’s Youth Opportunities Program (Blattman, Fiala and Martinez, 2014), it would have corroborated the reported randomization mechanism while flagging selection and missing-data risks that matter for interpretation and design; in a field experiment in Bangladesh, randomization integrity is itself questioned, in line with concerns raised by an independent team of investigators.
Remote audits are not a panacea: they detect only image‑aligned deviations, and a large pp‑value is not a certificate of perfection. Yet the audit’s low cost, finite‑sample validity, and preregistrability make it a practically useful tool when ground measurement is slow or difficult. Because the procedure can be run ex post as well as ex ante, it also enables a broader agenda: a grand audit of field‑experimental assignments using only archived pre‑treatment imagery and registered designs. As global archives deepen and off‑the‑shelf vision models improve (Li et al., 2020; Dosovitskiy et al., 2020; Liu et al., 2021), we recommend incorporating remote audits alongside conventional balance checks and process documentation in both preregistration and retrospective quality assurance. $\square$
@@
Proof. Condition on $(\boldsymbol{\phi},H)$. Under the null, $A\sim\Omega$ and $A^{(1)},\ldots,A^{(B)}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\Omega$ are exchangeable. Applying the fixed, measurable map $g(\cdot)$ to each assignment yields exchangeable statistics $\big(T,T^{(1)},\ldots,T^{(B)}\big)$. Therefore the rank of $T$ among these B+1B{+}1 values is uniformly distributed on $\{1,\ldots,B{+}1\}$ (with ties handled by the $\geq$ rule or broken at random), which implies that $p$ is (super-)uniform and $\Pr(p\leq\alpha\mid\boldsymbol{\phi},H)\leq\alpha$. When $\Omega$ imposes block or cluster totals, exchangeability holds conditional on those totals, so the same argument applies. $\square$
Remark. Refitting the predictive model within each resample is what ensures that $g(\boldsymbol{\phi},\cdot,H)$ treats every draw from $\Omega$ symmetrically; reusing a learner trained only on the realized AA can break exchangeability. See Candes et al. (2018) and Hennessy et al. (2016) for more information.